OpenStudy (anonymous):

Solve the system of inequalities by graphing. y≤-x-2 y>3x-3

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

@satellite73

OpenStudy (mathstudent55):

Do you know how to graph the equation of a line? For example, can you graph y = x - 2?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Ok. To graph the inequalities you have, you start with the corresponding line. Let's do it one inequality at a time.

OpenStudy (anonymous):

think it is actually \(y=-x-2\) right?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

now i will be quiet so @mathstudent55 will explain

OpenStudy (anonymous):

no

OpenStudy (anonymous):

xD

OpenStudy (mathstudent55):

@satellite73 Thanks. Let's start with the first inequality. First think of its corresponding equation. All you do is change the inequality sign to an equal sign. y = -2x - 2

OpenStudy (mathstudent55):

Since the original inequality has a less than or EQUAL to sign, graph the corresponding equation with a solid line.

OpenStudy (anonymous):

????

OpenStudy (mathstudent55):

@Hero that has been taken care of above. See satellite73's comment. Then take one point not on the line and test it in the inequality

OpenStudy (mathstudent55):

Oh, now I get it. thanks @Hero

OpenStudy (anonymous):

which one do i graph

OpenStudy (mathstudent55):

Ok, sorry about that. The first inequality is y ≤ - x - 2 Let's graph its corresponding equation using a solid line. y = -x - 2

OpenStudy (anonymous):

so graph y=-x - 2

OpenStudy (mathstudent55):

Yes. 

OpenStudy (anonymous):

kk

OpenStudy (mathstudent55):